Use the sign of the derivative to establish the inequaltiy ln(1 + x) > x − x 2 2 , ∀ x > 0


1
Expert's answer
2021-05-10T14:22:04-0400

Define function "f(x)" as:

"f(x)=\frac{x^2}{2}+ln(1+x)-x"


Note that: "f(0)=0"


Let us find first derivative of function "f(x)" :

"f'(x)=\frac{x^2}{x+1}\ ,"


Note that "f'(x)>0" for all "x>0"


Since "f(0)=0" and "f'(x)>0" for all "x>0" it follows that:


"\frac{x^2}{2}+ln(1+x)-x>0\Rightarrow ln(1+x)>x-\frac{x^2}2{}" , for all "x>0"


Hence Proved.